# Pinhole Effect on Linkability and Dispersion in Speaker Anonymization

### v1

@@ IV-C Performance metrics
Table I shows the baseline performance of the two anonymization systems, SYS1 and SYS2, using the any-to-one mapping to a universal voice. The ASV evaluations were conducted in a a gender-dependent manner. The average EERs across the four subsets are included. Comparing to the EERs in the ORG column, when no anonymization was applied, both anonymization systems increase tremendously the EER to over $400\%$. Both anonymization systems obtained WERs comparable with the original speech (ORG) without anonymization. These results show that the anonymization systems achieve good privacy preservation while retaining the linguistic information for downstream tasks.
@@ V Conclusion
We have introduced the pinhole effect as a conceptual framework to explain the identity linkability behavior frequently observed in speaker anonymization systems. By modeling the anonymization process as a mapping function from original speaker identities to pseudo speakers, we examined two key strategies: mapping to a common pseudo speaker (any-to-one) and mapping to distinct pseudo speakers (any-to-any). Our analysis shows that anonymizing each utterance to a unique pseudo speaker significantly reduces speaker linkability by increasing the dispersion in the anonymized speaker space. While both mapping strategies achieve comparable de-identification performance (i.e, the anonymized speech cannot be reliably traced back to the original speaker), the use of unique pseudo speaker offers a clear advantage in lowering linkability, which is a desirable property in privacy-preserving speech processing.

### v2

@@ IV-C Performance metrics
Table I shows the baseline performance of the two anonymization systems, SYS1 and SYS2, using the any-to-one mapping to a universal voice. The ASV evaluations were conducted in a a gender-dependent manner. The average EERs across the four subsets are included. Comparing to the EERs in the ORG column, when no anonymization was applied, both anonymization systems increase tremendously the EER to over $30\%$. Both anonymization systems obtained WERs comparable with the original speech (ORG) without anonymization. These results show that the anonymization systems achieve good privacy preservation while retaining the linguistic information for downstream tasks.
@@ F M F M
IV-E Statistical significance test
We conducted bootstrap resampling to estimate the 9595% confidence intervals (CIs) of the EER differences between a2a and a2o, denoted as $d_{\text{EER}}$, for each test condition in Table III. The detailed procedure is described in the Appendix. As shown in Table V, the null hypothesis value (zero difference) falls outside the 9595% CI, indicating that the EER differences are statistically significant ($p<0.05$).
TABLE V: Bootstrap-estimated 95% confidence intervals (CIs) of EER differences ($d_{\text{EER}}$) for four test conditions.
System Libri-Dev (F) Libri-Dev (M) Libri-Test (F) Libri-Test (M)
SYS1 (Mean) 1.75 3.17 1.77 2.55
95% CI 0.27–2.61 0.15–6.05 0.45–4.31 0.52–6.50
SYS2 (Mean) 2.28 3.42 2.11 4.36
95% CI 0.89–5.57 0.77–6.41 0.54–4.47 1.12–7.23
Similarly, Table VI shows the bootstrap confidence intervals computed for each test condition in Table IV. In this case, the 9595% CI includes the null hypothesis value, indicating that there is no substantial difference ($p>0.05$). Any-to-one and any-to-any mappings achieve a comparable level of de-identification performance, consistent with the pinhole effect.
TABLE VI: Bootstrap-estimated 95% confidence intervals (CIs) of EER differences ($d_{\text{EER}}$) for de-identification performance.
System Libri-Dev (F) Libri-Dev (M) Libri-Test (F) Libri-Test (M)
SYS1 (Mean) 0.41 -0.38 -0.89 2.43
95% CI -1.34–2.74 -2.84–2.69 -4.04–3.34 -1.34–6.44
SYS2 (Mean) 0.36 -0.14 1.54 0.61
95% CI -2.78–3.67 -2.05–1.68 -2.74–5.81 -2.36–4.71
@@ V Conclusion
We have introduced the pinhole effect as a conceptual framework to explain the identity linkability behavior frequently observed in speaker anonymization systems. By modeling the anonymization process as a mapping function from original speaker identities to pseudo speakers, we examined two key strategies: mapping to a common pseudo speaker (any-to-one) and mapping to distinct pseudo speakers (any-to-any). Our analysis shows that anonymizing each utterance to a unique pseudo speaker significantly reduces speaker linkability by increasing the dispersion in the anonymized speaker space. While both mapping strategies achieve comparable de-identification performance (i.e, the anonymized speech cannot be reliably traced back to the original speaker), the use of unique pseudo speaker offers a clear advantage in lowering linkability, which is a desirable property in privacy-preserving speech processing. [Statistical Significance Test Using Bootstrap] To assess whether the EER difference between the two methods (a2a and a2o) is statistically significant, a bootstrap procedure was employed as follows.
  1. 1.
    Generate Bootstrap Samples: From the original ASV score set of size $n$, multiple bootstrap samples were created by resampling with replacement. Each bootstrap sample contained $n$ data points, allowing duplicates; hence, some scores from the original dataset could appear multiple times, while others might not appear at all.
  2. 2.
    Compute the Statistic: For each bootstrap sample, the equal error rates (EERs) of both methods (a2a and a2o) were computed, and their difference was calculated as
    $d_{\text{EER}}=\text{EER}_{\text{a2a}}-\text{EER}_{\text{a2o}}.$
  3. 3.
    Build the Bootstrap Distribution: The resampling and computation process was repeated $1000$ times to obtain a bootstrap distribution of $d_{\text{EER}}$.
  4. 4.
    Determine the Confidence Interval: The bootstrap differences were sorted, and the $2.5$th and $97.5$th percentiles were taken to form the 9595% confidence interval (CI).
  5. 5.
    Test for Statistical Significance: If the null hypothesis value (zero difference) lay outside the 9595% CI, the null hypothesis was rejected, indicating that the EER difference between the two methods was statistically significant.